# Explainable Anomaly Detection in Images and Videos: A Survey

### v2

@@ Abstract
Anomaly detection and localization of visual data, including images and videos, are of great significance in both machine learning academia and applied real-world scenarios. Despite the rapid development of visual anomaly detection techniques in recent years, the interpretations of these black-box models and reasonable explanations of why anomalies can be distinguished out are scarce. This paper provides the first survey concentrated on explainable visual anomaly detection methods. We first introduce the basic background of image-level and video-level anomaly detection. Then, as the main content of this survey, a comprehensive and exhaustive literature review of explainable anomaly detection methods for both images and videos is presented. Next, we analyze why some explainable anomaly detection methods can be applied to both images and videos and why others can be only applied to one modality. Additionally, we provide summaries of current 2D visual anomaly detection datasets and evaluation metrics. Finally, we discuss several promising future directions and open problems to explore the explainability of 2D visual anomaly detection. The related resource collection is given at https://github.com/wyzjack/Awesome-XAD.
@@
Anomaly Detection, Explainable Machine Learning, Image Processing, Video Processing.
@@ I Introduction
Anomaly detection plays a crucial role in society. Nowadays, with the development technique of computer vision, many researchers try to address 2D anomaly detection tasks, including image anomaly detection [1, 2, 3, 4] and video anomaly detection [5, 6, 7, 8, 9, 10, 11]. 2D anomaly detection aims to identify unusual or abnormal patterns in visual data, such as images and videos. It is used in various applications, such as surveillance, security, medical imaging, and industrial inspection. The basic idea behind 2D anomaly detection is to train a model to recognize normal patterns in a set of training data, and then use that model to identify instances where new, unseen data deviates from those normal patterns. Considering the nature of the data, most of the statistic methods, like isolation forest and KNN, that work well on tabular data cannot directly apply to image data. Deep learning-based methods are extensively employed in 2D anomaly detection. For example, Autoencoders and Variational Autoencoders (VAEs) are common architectures used in 2D anomaly detection. Autoencoders are neural networks that are trained to reconstruct their input by learning a compact representation of the data. VAEs are a probabilistic extension of autoencoders that learn a generative model of the data. Through training these models on a dataset of normal images, it can be used to identify images that deviate significantly from the normal patterns.
Although existing work has proved the potential of deep learning models to identify anomalies in visual data, the explainability of those works is limited. Explainability is defined as the ability to explain or provide meaning to humans in understandable terms. As for the 2D anomaly detection task, an explainable model should give appropriate and faithful reasons for making such decisions. In fact, model explainability represents one of the main issues concerning the adoption of data-driven algorithms in industrial environments [12, 13]. More importantly, providing explanations for AI systems is now both an ethical and a regulatory requirement, especially in fields where human safety is a concern. For example, in the biomedical image anomaly detection task, a deep learning model can give a precise prediction for X-ray images with respect to disease, but without any explanation. However, it is very risky to fully trust it. In many other cases, our goal is not only to detect anomalies but also to localize them and explain them so that we can troubleshoot those anomalies and fix them eventually. Thus, explainable 2D anomaly detection is necessary and important for a wide variety of human activities.
To tackle this issue, some methods have been proposed in recent years. For different tasks and user orientations, the explanation proposals are largely different, which causes the appearance of numerous explainable methods. For example, in the surveillance video anomaly detection task, [14] uses an autoencoder-based method to detect anomalies, and the prediction error is constructed as a saliency map with which off-the-shelf object detection and action recognition model are applied to explain anomalies post hoc. In this survey, we focus on existing methods related to 2D anomaly detection tasks. In particular, we classify them into two types: image anomaly detection and video anomaly detection.
To the best of our knowledge, anomaly detection has been widely studied, and there exists a lot of surveys [15, 16, 17, 18] about them. In contrast, there are a limited number of existing surveys on the topic of explainable anomaly detection [13]. Moreover, those existing surveys try to review all explanation methods with respect to all kinds of data types, which provide a good overview but lack insightful direction in 2D anomaly detection. In addition, there is no such survey in the explainable 2D anomaly detection area, and thus it is necessary to characterize the increasingly rich set of explainable 2D anomaly detection techniques available in the literature. In order to tackle this challenge, we conduct a comprehensive and organized survey of the current state-of-the-art techniques for explainable 2D anomaly detection. By providing a refined taxonomy that accommodates the growing variety of techniques being developed, and realistic and inspiring future directions, this survey intends to give readers a thorough understanding of the various methods proposed for explainable 2D anomaly detection, including their advantages and disadvantages, and to assist them in finding the explainable anomaly detection technique most suited to their needs or a research topic they are interested in.
@@
  • We collect and summarize explainable methods in 2D anomaly detection tasks, and propose the taxonomy of these explainable methods.
  • We thoroughly compare these methods, outline their advantages and disadvantages, and suggest potential directions to address existing drawbacks.
  • We fill in the gaps for current surveys that give a brief introduction to the whole anomaly detection tasks but ignore the particularity of explainable methods in the 2D anomaly detection task.
@@ II Background
is of considerable significance in both academia and industry, with wide-range application scenarios including industrial inspection, medical health, and website management, etc. Image anomaly detection can be broadly classified into two sub-tasks: semantic anomaly detection [19, 20, 21, 1] and sensory anomaly detection in images [22, 23, 3] and some approaches can possess both functions. For one-class classification, classification datasets, such as CIFAR-10 and CIFAR-100 [24], are usually employed to create a one-class or leave-one-out setting or dataset-wise out-of-distribution setting. [21] put forward a distribution-shifting transformations-based contrastive learning scheme. [20] seamlessly combine feature-level self-supervised learning with ad hoc feature adversarial perturbation for state-of-the-art performance. For both image-level anomaly detection and localization, the most commonly used dataset is MVTec AD [25][22] pioneers image-level anomaly detection and segmentation through the disparity in output between the teacher and student models. More recently, [3] exploits an ImageNet-pretrained network for patch-level feature embedding extraction and proposes an effective core-set selection and representation distance calculation for anomaly score design, exhibiting the best performance so far.
@@
Image Attention-based 2020 GradCon [35] C10, MN, fMN
@@
[…]
@@
2022 Diffusion-anomaly[47] BRATS, CheXpert
@@
[…]
@@
Image+Video Attention-based 2020 AD-FactorVAE [58] MN, MVTAD, Ped
@@
Video Attention-based 2020 Self-trained DOR [61] Ped2, Subway, UMN
@@
[…]
@@
III Explainable 2D Anomaly Detection
Explainable AI is essential to improving trust and transparency of AI-based systems  [71]. In anomaly detection, explainability is an ethical and regulatory requirement particularly in safety-critical domains [13, 16]. We list the state-of-the-art explainable anomaly detection methods in images and videos in Tab. I.
@@ III-A Explainable Image Anomaly Detection
Figure 1: Image from [23]. The proposed CAVGA approach employs the complementary guided attention loss to force the attention map to cover all the normal regions and meanwhile suppress the anomalous attention on the normal images.
For image anomaly detection, either semantic anomaly detection or pixel-level anomaly segmentation, explainability, and interpretability are of great significance to the understanding of why specific samples are considered anomalies and the development of more advanced approaches. Explainable image anomaly detection techniques can be broadly grouped into three kinds: Attention-based IAD, Input-perturbation-based IAD, and Generative-model-based IAD. The empirical result comparison on the public IAD benchmark MVTec AD is provided in Tab. II, where the Generative-model-based methods achieve the best performance.
TABLE II: Pixel-wise category-specific AUROC and mean AUROC for the MVTec-AD dataset [25] of explainable IAD methods.
Attention-based Generative-model-based
AD-FactorVAE [58] CAVGA [23] FCDD [36] SSPCAB [34] FastFlow [45] CFLOW-AD [60]
Carpet 0.78 - 0.99 0.950 0.994 0.9925
Grid 0.73 - 0.95 0.995 0.983 0.9899
Leather 0.95 - 0.99 0.995 0.995 0.9966
Tile 0.80 - 0.98 0.993 0.963 0.9801
Wood 0.77 - 0.94 0.968 0.970 0.9665
Bottle 0.87 - 0.96 0.988 0.977 0.9898
Cable 0.90 - 0.93 0.960 0.984 0.9764
Capsule 0.74 - 0.95 0.931 0.991 0.9898
Hazelnut 0.98 - 0.97 0.998 0.991 0.9889
Metal Nut 0.94 - 0.98 0.989 0.985 0.9856
Pill 0.83 - 0.97 0.975 0.992 0.9895
Screw 0.97 - 0.93 0.998 0.994 0.9886
Toothbrush 0.94 - 0.95 0.981 0.989 0.9893
Transistor 0.93 - 0.90 0.870 0.973 0.9799
Zipper 0.78 - 0.98 0.990 0.987 0.9908
Mean 0.86 0.93 0.96 0.972 0.985 0.9862
Attention-based IAD
@@
Input-perturbation-based IAD
TABLE III: Summary of the detailed perturbation formula in input-perturbation-based image anomaly detection task. $\epsilon$ denotes the perturbation magnitude, sign denotes the signum function, $S(x)$ is some score function of $x$ and $D_{\text{in}}^{\text{val}}$ denotes the in-distribution validation set. The dashed line means that the first three methods utilize pre-trained classifier output for $S(x)$ while the last one uses a self-supervised learned classifier.
@@
ODIN (2017) $-\epsilon\text{sign}(-\nabla_{x}\log S(x))$ searched maximum softmax probability
Mahalanobis (2018) $+\epsilon\text{sign}(\nabla_{x}S(x))$ searched Mahalanobis distance-based confidence score
Generalized ODIN (2020) $-\epsilon\text{sign}(-\nabla_{x}S(x))$ $\mathop{\mathrm{argmax}}_{\epsilon}\sum_{x\in D_{\text{in}}^{\text{val}}}S(x)$ maximum softmax probability
[…]
@@
where $x$ denotes the input data or feature, and $x^{\prime}$ is the perturbed data or feature. The perturbation term is usually in the form of a perturbation magnitude multiplying the gradient of some score function w.r.t. the input. Then the perturbed data or feature serves as input for anomaly score calculation.  [37] is the pioneering work adopting input perturbation technique in anomaly detection. The proposed ODIN method takes the gradient of a pretrained network and aims to increase the softmax score of any given input. [39] generalizes ODIN via finding the maximum magnitude which maximizes the sum of the scores w.r.t. the in-distribution validation dataset. The generalized ODIN approach further enhances detection performance consistently. Instead of directly using softmax output scores,  [38] proposes Mahalanobis distance-based confidence score, and the perturbation is introduced to enhance the confidence score. Different from these OOD detection work which computes the gradient using a pretrained network, [20] trains a neural network with feature-level self-supervised learning and adds perturbation on the feature embeddings using the trained classifier network. The detailed comparison of the input perturbations of these literature is summarized in Tab. III. Despite the differences among these input-perturbation methods, the rationale that they can explain anomalous patterns is the same: due to the characteristics of the anomaly detection problem anomalies are lower in quantity compared to the normal samples and the modes of anomalies are more diverse when applying the same amount of perturbation, the score alternations of them will be different from those of the normal data. Therefore we are able to examine the effect of output via input changes to generate reasonable explanations.
Generative-model-based IAD
TABLE IV: Summary of the generative-model-based image anomaly detection methods. Here CVPRW denotes CVPR Workshop. The abbreviations recon.error represents reconstruction error, disc.score represents discrimination score and recon.feature symbols reconstruction features respectively.
Method Venue Generative model type Anomaly score function
AnoGAN [40] IPMI 2017 Generative Adversarial Network recon.error $+$ disc.score
ALAD [41] ICDM 2018 Generative Adversarial Network recon.error $+$ disc.score
f-AnoGAN [42] MIA 2019 Generative Adversarial Network recon.error
LSAND [59] CVPR 2019 Autoregressive Model recon.error $+$ likelihood
Genomics-OOD [43] NeurIPS 2019 Autoregressive Model likelihood ratio
AD-FactorVAE [58] CVPR 2020 Variational Autoencoder gradient-based attention map
Likelihood Regret [44] NeurIPS 2020 Variational Autoencoder likelihood regret
FastFlow [45] arXiv 2021 Normalizing Flow likelihood
CFLOW-AD [60] WACV 2022 Normalizing Flow likelihood
AnoDDPM [46] CVPRW 2022 Diffusion Model recon.error
Diffusion-anomaly [47] MICCAI 2022 Diffusion Model recon.error
DDPM [48] MICCAI 2022 Diffusion Model recon.error
DiAD [48] AAAI 2024 Diffusion Model similarity
@@
Generative models (GM) are widely adopted in 2D anomaly detection. As summarized in Tab. IV, the most commonly used GMs include Generative Adversarial Networks (GAN), Autoregressive Models, Variational Autoencoders (VAE), Normalizing Flows, Diffusion Models and other variants of autoencoders. [40] is the first work to seamless combine GAN with image anomaly detection and localization task, where the reconstruction error of GAN and the discrimination score from the discriminator are jointly used for anomaly score design. This is because the GAN fails to reconstruct the anomalies well and the discriminator will regard the anomalies as fake since they only encounter normal ones in the training phase. [59] employs a deep autoencoder together with an autoregressive model for density estimation which learns the distribution of the latent representations via maximum likelihood principles and proves empirically this is an efficient regularizer for anomaly detection. Furthermore, [43] demonstrates that likelihood from generative model-based anomaly detection methods is problematic due to the background statistics and proposes a novel likelihood ratio-based anomaly score. The key finding is that likelihood value is mainly controlled by the background pixels while likelihood ratios concentrate more on the semantic information, which explains why likelihood ratio is a more preferable choice for anomaly score for one-class classification. The formulation of the proposed likelihood ratio is
@@
where $x$ is the data sample, $p_{\theta}(\cdot)$ is the model trained employing in-distribution data and $p_{\theta_{0}}(\cdot)$ is a background model which captures general background statistics. Similarly, [44] addresses the problem that generative models, especially VAEs, are inclined to assign higher likelihoods on some types of OOD samples through defining a new OOD detection score dubbed as Likelihood Regret. They show that the Likelihood regret can be viewed as a specific form of log ratio of the likelihood from VAE. It is effective in separating anomalies out in that the increase of likelihood by substituting the model configuration with the optimal one for the single sample ought to be comparatively small for normal samples and large for anomalous ones thanks to the fact that the VAE is trained well on the normal training data.
@@
represents the evidence lower bound, (ELBO), $\phi$ and $\theta$ are the parameters of the encoder and decoder respectively, $\phi^{*}$ and $\theta^{*}$ are the network parameters maximizing the population log-likelihood over the training dataset, and $\widehat{\tau}(x)$ denotes the optimal parameter maximizing the ELBO of the test input $x$.  [58] also utilizes VAE for anomaly detection, but instead uses a gradient-based attention map generated from the latent space of VAE for anomaly detection and localization in images. FastFlow [45] employs 2D normalizing flows for efficient and unsupervised anomaly detection and localization, providing explainability through intuitive visualizations of the transformed features for anomaly identification. CFLOW-AD [60] leverages conditional normalizing flows for unsupervised anomaly detection, offering explainability through the generative model’s ability to localize anomalies by contrasting reconstructed normal and anomalous regions, thereby enhancing the interpretability of detection results. More recently, diffusion models exhibit superior image generation ability and hence emerge in anomaly detection research. For example, AnoDDPM [46] proposes a partial diffusion strategy with simplex noise for medical image anomaly detection and localization. [47] integrates DDPM [76] with a binary classifier for medical image anomaly segmentation and [48] devises a diffusion model for fast brain image anomaly segmentation. More recently, DiAD [49] develops a framework for detecting anomalies across multiple classes, integrating an autoencoder in pixel space, a semantic-guided network in latent space connected to the denoising network of the diffusion model, and a pre-trained extractor in the feature space, with innovations like spatial-aware feature fusion for enhanced reconstruction accuracy and semantic preservation. The critical advantage of the Diffusion Model for explainability is that each step in the forward step and reverse step has a closed from, which enables people to visualize the latent outputs of each step or layer in order to investigate how anomalies are ruled out in the diffusion process and where the model pay more attention.
@@
Foundation-model-based IAD

### v4

# Unveiling the Unseen: A Comprehensive Survey on Explainable Anomaly Detection in Images and Videos

@@ Abstract
Anomaly detection and localization in visual data, including images and videos, are crucial in machine learning and real-world applications. Despite rapid advancements in visual anomaly detection (VAD), interpreting these often black-box models and explaining why specific instances are flagged as anomalous remains challenging. This paper provides the first comprehensive survey focused specifically on explainable 2D visual anomaly detection (X-VAD), covering methods for both images (IAD) and videos (VAD). We first introduce the background of IAD and VAD. Then, as the core contribution, we present a thorough literature review of explainable methods, categorized by their underlying techniques (e.g., attention-based, generative model-based, reasoning-based, foundation model-based). We analyze the commonalities and differences in applying these methods across image and video modalities, highlighting modality-specific challenges and opportunities for explainability. Additionally, we summarize relevant datasets and evaluation metrics, discussing both standard performance metrics and emerging approaches for assessing explanation quality (e.g., faithfulness, stability). Finally, we discuss promising future directions and open problems, including quantifying explanation quality, explaining diverse AD paradigms (SSL, zero-shot), enhancing context-awareness, leveraging foundation models responsibly, and addressing real-world constraints like efficiency and robustness. A curated collection of related resources is available at https://github.com/wyzjack/Awesome-XAD.
@@
Anomaly Detection, Explainable Machine Learning, Image Processing, Video Processing, Computer Vision.
@@ I Introduction
Anomaly detection plays a crucial role in society. Nowadays, with the development technique of computer vision, many researchers try to address visual anomaly detection tasks, specifically focusing on 2D data which primarily includes images and videos (referred to as 2D visual anomaly detection hereafter). Relevant tasks include image anomaly detection (IAD) [1, 2, 3, 4] and video anomaly detection (VAD) [5, 6, 7, 8, 9, 10, 11] (3D point cloud anomaly detection works [12, 13, 14, 15] are out of the scope of this paper). 2D visual anomaly detection aims to identify unusual or abnormal patterns within these visual modalities. It finds applications in various domains, such as surveillance, security, medical imaging, and industrial inspection. The fundamental idea is to train a model on normal patterns and then identify deviations in new, unseen data. Due to the high-dimensional nature of visual data, traditional statistical methods effective for tabular data often fall short. Consequently, deep learning-based methods are extensively employed. Common architectures like Autoencoders (AEs) and Variational Autoencoders (VAEs) learn compact representations or generative models of normal data, enabling the identification of outliers based on reconstruction errors or likelihood scores.
While deep learning models show promise in identifying visual anomalies, their inherent ”black-box” nature often limits their explainability. Explainability, in this context, refers to the ability of a model to provide understandable reasons for its decisions to human users [16, 17]. For 2D visual anomaly detection, an explainable model should offer clear and faithful justifications for why a specific image region or video segment is flagged as anomalous. The lack of transparency hinders trust and practical adoption, particularly in safety-critical domains [18, 19]. For instance, in medical diagnosis, simply predicting a disease from an X-ray without explaining which features led to the conclusion is insufficient for clinical acceptance. Providing explanations is becoming an ethical and regulatory necessity [20]. Furthermore, understanding why something is anomalous is often crucial for troubleshooting and remediation in applications like industrial quality control. Therefore, developing explainable 2D visual anomaly detection methods is essential.
To address the need for explainability, various methods have emerged recently. These approaches differ significantly based on the specific task, data modality (image vs. video), and the target audience for the explanation. For example, in surveillance video anomaly detection, [21] employed an autoencoder and used the reconstruction error map (saliency map) combined with off-the-shelf object/action recognition models to provide post-hoc explanations. This survey focuses specifically on explainable methods developed for or applied to 2D visual anomaly detection tasks, categorizing them primarily based on the underlying technique used to achieve explainability within the context of IAD and VAD.
Anomaly detection in general is a well-studied field with numerous surveys [22, 23, 24, 25, 26]. Explainable AI (XAI) also has comprehensive reviews covering various methods and metrics [27, 17, 28, 16, 29, 20, 30]. However, surveys specifically focusing on explainable anomaly detection are fewer [19, 31]. Notably, the recent survey by Li et al. [32] provides a valuable overview of explainable anomaly detection across different data types. While insightful, its broad scope means it provides less depth on the unique challenges and techniques specific to the visual domain (images and videos). Our survey aims to fill this gap by concentrating exclusively on explainable 2D visual anomaly detection. We delve deeper into the specific explainability methods tailored for or applied to images and videos, analyze the commonalities and differences between IAD and VAD explainability, discuss domain-specific datasets and evaluation metrics relevant to explainability, and propose future directions pertinent to visual data. Our literature search aimed for comprehensive coverage of papers explicitly addressing explainability in visual anomaly detection, retrieved through systematic searches of major computer vision and machine learning venues and citation chasing. This focused approach allows for a more detailed characterization of the state-of-the-art in this rapidly evolving subfield, offering a refined taxonomy and actionable insights for researchers and practitioners working with visual data.
@@
  • We provide the first focused survey on explainable methods specifically for image and video anomaly detection.
  • We propose a taxonomy tailored to these explainable visual methods, comparing them thoroughly, outlining their advantages, disadvantages, and specific applicability to IAD versus VAD.
  • We analyze the unique aspects of explainability in visual domain, including relevant datasets and evaluation considerations beyond standard anomaly detection metrics.
  • We identify and discuss potential future research directions specific to explainable 2D visual anomaly detection.
@@ II Background
is of considerable significance in both academia and industry, with wide-range application scenarios including industrial inspection, medical health, and website management, etc. Image anomaly detection can be broadly classified into two sub-tasks: semantic anomaly detection [33, 34, 35, 1] and sensory anomaly detection in images [36, 37, 3] and some approaches can possess both functions. For one-class classification, classification datasets, such as CIFAR-10 and CIFAR-100 [38], are usually employed to create a one-class or leave-one-out setting or dataset-wise out-of-distribution setting. [35] put forward a distribution-shifting transformations-based contrastive learning scheme. [34] seamlessly combine feature-level self-supervised learning with ad hoc feature adversarial perturbation for state-of-the-art performance. For both image-level anomaly detection and localization, the most commonly used dataset is MVTec AD [39][36] pioneers image-level anomaly detection and segmentation through the disparity in output between the teacher and student models. More recently, [3] exploits an ImageNet-pretrained network for patch-level feature embedding extraction and proposes an effective core-set selection and representation distance calculation for anomaly score design, exhibiting the state-of-the-art performance so far.
@@
Image Attention 2020 GradCon [49] C10, MN, fMN
@@
Input perturbation 2018 ODIN [51] TinyImageNet, LSUN, iSUN
@@
2022 Diffusion-anomaly [61] BRATS
@@
2025 DeCo-Diff [64] MVTAD, VisA
Foundation model 2023 WinCLIP [65] MVTAD, VisA
@@
2024 CLIP-SAM [73] MVTAD, VisA
2024 AnomalyClip [74] MVTAD, VisA
2024 AA-CLIP [75] MVTAD, VisA, MPDD, BTAD, Retina OCT
2024 MVFA [76] BrainMRI
2024 InCTRL [77] MVTAD, VisA, BTAD MPDD, BrainMRI
2025 Bayes-PFL [78] MVTAD, VisA, KSDD2, BTAD, DAGM, BrainMRI
2025 LogSAD [79] MVTAD, VisA, MVTLC
2025 LogicAD [80] MVTLC
2025 Anomaly-OV [81] MVTAD, VisA, BTAD, MPDD, BrainMRI
2025 UniVAD [81] MVTAD, VisA, MVTLC, BrainMRI
Image+Video Attention 2020 AD-FactorVAE [82] MN, MVTAD, Ped
@@
Video Attention 2020 Self-trained DOR [85] Ped2, Subway, UMN
@@
2024 VERA [91] UCF, XD
2024 AnomalyRuler [92] Ped2, Ave, STC, UBnormal
2024 LAVAD [93] UCF, XD
2024 Holmes-VAU [94] UCF, XD
2025 Anomize [95] UCF, XD
Intrinsic interpretable 2017 JDR [96] Ped2, Ave
@@
Memory 2020 Memory AD [100] Ped2, Ave, STC
2021 DLAN-AC [101] Ped2, Ave, STC
III Explainable 2D Visual Anomaly Detection
Explainable AI is essential for improving trust and transparency in AI-based systems [102]. In anomaly detection, explainability is an ethical and regulatory requirement, particularly in safety-critical domains [19, 24]. We provide a detailed overview of methods that incorporate explainability into 2D visual anomaly detection, categorized by their primary approach and modality.
@@ III-A Explainable Image Anomaly Detection
Figure 1: [37]. The proposed CAVGA approach employs the complementary guided attention loss to force the attention map to cover all the normal regions and meanwhile suppress the anomalous attention on the normal images.
For image anomaly detection, either semantic anomaly detection or pixel-level anomaly segmentation, explainability, and interpretability are of great significance to the understanding of why specific samples are considered anomalies and the development of more advanced approaches. Explainable image anomaly detection techniques can be broadly grouped into three kinds: Attention-based IAD, Input-perturbation-based IAD, and Generative-model-based IAD.
III-A1 Attention-based IAD
@@
III-A2 Input-perturbation-based IAD
TABLE II: Summary of the detailed perturbation formula in input-perturbation-based image anomaly detection task. $\epsilon$ denotes the perturbation magnitude, sign denotes the signum function, $S(x)$ is some score function of $x$ and $D_{\text{in}}^{\text{val}}$ denotes the in-distribution validation set. The dashed line means that the first three methods utilize pre-trained classifier output for $S(x)$ while the last one uses a self-supervised learned classifier.
@@ III-A2 Input-perturbation-based IAD
ODIN [51] $\epsilon\text{sign}(-\nabla_{x}\log S(x))$ searched maximum softmax probability
Mahalanobis [52] $\epsilon\text{sign}(\nabla_{x}S(x))$ searched Mahalanobis distance-based confidence score
Generalized ODIN [53] $\epsilon\text{sign}(-\nabla_{x}S(x))$ $\mathop{\mathrm{argmax}}_{\epsilon}\sum_{x\in D_{\text{in}}^{\text{val}}}S(x)$ maximum softmax probability
SLA2[34] $\epsilon\nabla_{x}\log S(x)$ searched maximum softmax probability
@@
where $x$ denotes the input data or feature, and $x^{\prime}$ is the perturbed data or feature. The perturbation term is usually in the form of a perturbation magnitude multiplying the gradient of some score function w.r.t. the input. Then the perturbed data or feature serves as input for anomaly score calculation. [51] is the pioneering work adopting input perturbation technique in anomaly detection. The proposed ODIN method takes the gradient of a pretrained network and aims to increase the softmax score of any given input. [53] generalizes ODIN via finding the maximum magnitude which maximizes the sum of the scores w.r.t. the in-distribution validation dataset. The generalized ODIN approach further enhances detection performance consistently. Instead of directly using softmax output scores, [52] proposes Mahalanobis distance-based confidence score, and the perturbation is introduced to enhance the confidence score. Different from these OOD detection work which computes the gradient using a pretrained network, [34] trains a neural network with feature-level self-supervised learning and adds perturbation on the feature embeddings using the trained classifier network. The detailed comparison of the input perturbations of these literature is summarized in Tab. II. Despite the differences among these input-perturbation methods, the rationale that they can explain anomalous patterns is the same: due to the characteristics of the anomaly detection problem anomalies are lower in quantity compared to the normal samples and the modes of anomalies are more diverse when applying the same amount of perturbation, the score alternations of them will be different from those of the normal data. Therefore we are able to examine the effect of output via input changes to generate reasonable explanations.
III-A3 Generative-model-based IAD
@@
Generative models (GM) are widely adopted in 2D anomaly detection. The most commonly used GMs include Generative Adversarial Networks (GAN), Autoregressive Models, Variational Autoencoders (VAE), Normalizing Flows, Diffusion Models and other variants of autoencoders. [54] is the first work to seamless combine GAN with image anomaly detection and localization task, where the reconstruction error of GAN and the discrimination score from the discriminator are jointly used for anomaly score design. This is because the GAN fails to reconstruct the anomalies well and the discriminator will regard the anomalies as fake since they only encounter normal ones in the training phase. [83] employs a deep autoencoder together with an autoregressive model for density estimation which learns the distribution of the latent representations via maximum likelihood principles and proves empirically this is an efficient regularizer for anomaly detection. Furthermore, [57] demonstrates that likelihood from generative model-based anomaly detection methods is problematic due to the background statistics and proposes a novel likelihood ratio-based anomaly score. The key finding is that likelihood value is mainly controlled by the background pixels while likelihood ratios concentrate more on the semantic information, which explains why likelihood ratio is a more preferable choice for anomaly score for one-class classification. The formulation of the proposed likelihood ratio is
@@
where $x$ is the data sample, $p_{\theta}(\cdot)$ is the model trained employing in-distribution data and $p_{\theta_{0}}(\cdot)$ is a background model which captures general background statistics. Similarly, [58] addresses the problem that generative models, especially VAEs, are inclined to assign higher likelihoods on some types of OOD samples through defining a new OOD detection score dubbed as Likelihood Regret. They show that the Likelihood regret can be viewed as a specific form of log ratio of the likelihood from VAE. It is effective in separating anomalies out in that the increase of likelihood by substituting the model configuration with the optimal one for the single sample ought to be comparatively small for normal samples and large for anomalous ones thanks to the fact that the Variational AE is trained well on the normal training data.
@@
represents the evidence lower bound, (ELBO), $\phi$ and $\theta$ are the parameters of the encoder and decoder respectively, $\phi^{*}$ and $\theta^{*}$ are the network parameters maximizing the population log-likelihood over the training dataset, and $\widehat{\tau}(x)$ denotes the optimal parameter maximizing the ELBO of the test input $x$.  [82] also utilizes VAE for anomaly detection, but instead uses a gradient-based attention map generated from the latent space of VAE for anomaly detection and localization in images. FastFlow [59] employs 2D normalizing flows for efficient and unsupervised anomaly detection and localization, providing explainability through intuitive visualizations of the transformed features for anomaly identification. CFLOW-AD [84] leverages conditional normalizing flows for unsupervised anomaly detection, offering explainability through the generative model’s ability to localize anomalies by contrasting reconstructed normal and anomalous regions, thereby enhancing the interpretability of detection results. More recently, diffusion models exhibit superior image generation ability and hence emerge in anomaly detection research. For example, AnoDDPM [60] proposes a partial diffusion strategy with simplex noise for medical image anomaly detection and localization. [61] integrates DDPM [107] with a binary classifier for medical image anomaly segmentation and [62] devises a diffusion model for fast brain image anomaly segmentation. More recently, DiAD [63] develops a framework for detecting anomalies across multiple classes, integrating an autoencoder in pixel space, a semantic-guided network in latent space connected to the denoising network of the diffusion model, and a pre-trained extractor in the feature space, with innovations like spatial-aware feature fusion for enhanced reconstruction accuracy and semantic preservation. Further exploration involves reformulating the diffusion process itself to better correct deviations from normality for multi-class unsupervised anomaly detection [64]. The critical advantage of the Diffusion Model for explainability is that each step in the forward step and reverse step has a closed form, which enables people to visualize the latent outputs of each step or layer in order to investigate how anomalies are ruled out in the diffusion process and where the model pay more attention.
@@
III-A4 Foundation-model-based IAD